# Models Matter: Setting Accurate Privacy Expectations for Local and Central Differential Privacy

### v1

@@ 3.1. Initial Prototypes
We also draw inspiration from prior work on visualizations of DP (Bullek et al., 2017; Nanayakkara et al., 2022; Pearce and Jiang, 2020; Smart et al., 2022; Near and Darais, 2020; Nanayakkara et al., 2023) to design our own diagrams that highlight how DP protects or fails to protect against the disclosures listed in Table 1. We developed our diagrams through an iterative process. We discussed the accuracy and clarity of initial diagrams as a group, and based on the discussion, iterated on our designs. In the end, we developed four diagrams—two for the local model and two for the central model—with slight differences in iconography. After initial interviews, we added a third variation for both models that included a caption. All diagrams used a vertical line to depict the “privacy barrier,” as in (Near and Darais, 2020), and used icons—most selected from the Noun Project [^3] https://thenounproject.com—to represent the different kinds of disclosures. Instead of using an illustration of a database, as in (Near and Darais, 2020; Xiong et al., 2022), we use an icon of a filing cabinet to represent the collected data. Representative diagrams can be found in Appendix D.
Following guidance from prior work, we also developed privacy labels to clearly demonstrate which kinds of information disclosures DP can protect against. Each row corresponds to a specific information disclosure and clarifies whether protection is offered against said disclosure. We tested three different versions of the tables (six distinct tables in total, across the two models). One version of the table listed only the disclosures against which DP can protect. Thus for the local model, this table had five rows, whereas for the central model, this version had only one row. This table uses a red circle-backslash symbol to indicate that a particular disclosure is not permitted. The other two versions always included information about all five disclosures, but used different iconography to depict protection or lack thereof. Both of these versions incorporated lock icons to indicate when DP protected against a particular kind of disclosure. In one of these tables, we use a green lock icon—as recommended in prior work on connection security icons (Felt et al., 2016)—to indicate safety, whereas a red unlocked icon indicates disclosures against which DP does not protect. The other table is in black-and-white and uses the presence or absent of a lock icon to indicate (lack of) protection. Chrome previously used lock icons to indicate connection security, but has recently backed away from this choice due to concerns about overtrust; some Chrome users incorrectly assumed that a lock icon was a reflection on the safety of the website itself rather than the connection (Chromium Blog, 2023; Ma et al., 2019). Varying the use of icons allowed us to evaluate their appropriateness in a DP context. Appendix D includes representative versions of our original privacy labels.
3.2. Protocol
@@
Next, participants are informed that the non-profit plans to “use an extra layer of privacy protection in order to protect patients’ medical information.” Then, they are shown the first explanation of this privacy protection. After reading the explanation, the participants are asked to explain how patient data will be protected in their own words, as in (Golla et al., 2018). Next, they are asked how they feel about the explanation, how well they feel that they understand the privacy protection after reading the explanation, what concerns they would have about sharing their data, and what else they would like to know about how patient data will be protected, adapting questions from (Redmiles et al., 2017). If the design under discussion includes the use of color, they are also asked about these color choices. Finally, they are asked how the explanation could be improved.
@@ 3.3. Participant Recruitment
3.4. Analysis
The interviewer first transcribed and summarized all the interviews. Next, an interview from each condition was selected at random, forming a set of six interviews. The first two authors reviewed these six transcripts to develop a set of codes, organized into four distinct themes (Appendix A). They then shared this codebook with the research team and modified it based on the group’s feedback. Finally, the same two authors coded all 24 interviews together [^4] We do not calculate or report inter-rater reliability (IRR) for two reasons. One, while calculating IRR can be useful to establish agreement before researchers divide a corpus to code different subsets individually, in our case both researchers coded all of the data together. Two, we are not seeking to make quantitative claims about our codes (McDonald et al., 2019). with the updated codebook.
3.5. Findings
3.5.1. Effectiveness of Initial Designs
@@
Responses to the metaphors were mixed. Some participants appreciated the concision of the metaphors, while others wanted more details. For example, one participant criticized an explanation’s brevity, saying it is “a little bit simple and […] doesn’t go into too many details.” (P8) In contrast, a different participant complimented this very quality by describing an explanation as “reader-friendly, very concise” (P5). This tension between accuracy and thoroughness of explanations on the one hand, and simplicity on the other has also been reported in other domains, such as explainable machine learning (Abdul et al., 2020) and privacy policies (Gluck et al., 2016b).
@@
Despite the fact that the diagrams showed the hacker to the left of the privacy barrier, two of the four participants in this condition nevertheless explicitly stated that the privacy barrier would protect their data from hackers. Many people may expect PETs to protect against hackers and criminals, making the protection offered by central DP alone somewhat unintuitive (Smart et al., 2022). We dropped the diagram explanations due to the pervasive confusion expressed by participants.
Xiong et al. (Xiong et al., 2022) previously investigated the use of diagrams for explaining location privacy and found less than ideal levels of comprehension, particularly for the local model, though they speculate that data quality issues with Amazon Mechanical Turk may be to blame. Alternatively, it is possible that data flow diagrams inherently overemphasize processes at the expense of clearly enumerating implications.
@@
All of our explanations were designed to communicate the implications of DP rather than the details of how DP works. Prior work has shown that explaining the process of adding noise to data is not enough to help people understand the consequences data sharing (Xiong et al., 2020). Nevertheless, omitting any discussion of process seems to leave people unsatisfied and confused. Most participants had questions about how the data protection worked. Providing a detailed explanation of the mathematical and technical details of DP is likely to overwhelm most people, but people nevertheless do want some information about how DP works—finding the right balance may be challenging. This finding aligns with prior work on explaining encryption. While explanations of encryption focused on outcome lead to greater perceived security than explanations focused on process, hybrid explanations that incorporate information on both process and outcome lead to the greatest perceived security (Distler et al., 2020). Prior work on metaphors for DP also found that some participants were interested in understanding how DP works (Karegar et al., 2022).
@@
[skipabove=] Design Changes: Another text was added to provide context about how DP works; we adapted a state-of-the-art text explanation by Xiong et al. (Xiong et al., 2020), while aiming for improved readability by eliminating terms like “database” and “aggregated.” We anticipated that this additional information on process could complement our other explanations that focus on implications.
3.5.2. Mental Models
@@
Some participants—especially in the diagram and privacy label conditions—reasoned about DP through comparisons to other PETs. For example, one participant understood the privacy barrier as “some kind of firewall that keeps [their] privacy safe” (P21). Encryption in particular was mentioned frequently, perhaps because it is a particularly familiar and ubiquitous PET or perhaps because participants associated our lock icons with encryption (Felt et al., 2016; Herzberg and Leibowitz, 2016). One participant, who assumed that encryption was the technology being described, wanted to know “what type of encryption” (P23) was used. Prior work has also found associations between DP and encryption, and found that associations with encryption correspond to higher trust (Karegar et al., 2022). DP is distinct from encryption, so while it may be possible to leverage people’s knowledge about encryption to construct better explanations of DP, it is also likely that associations with encryption may lead to misconceptions.
@@
Once again, while this model oversimplifies DP, it shares key elements with the truth and thus is likely useful overall. However, it is important for people to understand that the “fake” data nonetheless reveals useful information about the overall distribution; therefore, DP does not necessarily offer protection against inferential privacy risks (Kifer and Machanavajjhala, 2011, 2014).
3.5.3. Validating Design Changes.
@@
4. Large-Scale Evaluation
@@
Respondents who pass the comprehension check are then asked whether they trust the non-profit to protect patient privacy (Xiong et al., 2020). Next they are asked two questions related to self-efficacy. Finally, they are asked whether they would be willing to share their information with the non-profit. An open-text box asks them to explain their decision.
Respondents then answer five true/false questions on privacy expectations, followed by the Likert-scale questions about understanding and thoroughness. They are also invited to share feedback on the explanations in a free-response text box. When answering the above questions, respondents have the option to reread the descriptions of the scenario and privacy protection at any time. Next, respondents are asked about their familiarity with various PETs, including DP and a non-existent technology (“deliquescent security”). If they indicate familiarity with some of the listed technologies, they are asked which of the technologies (if any) was described in the survey. In a free-response text box, they are asked to explain their reasoning. Finally, respondents answer questions about themselves. In addition to standard demographic questions (i.e., age, income, race, ethnicity, gender, education, job field), the survey also includes measures of internet skill (Hargittai and Hsieh, 2012). The full survey instrument is included in Appendix B.
Table 2. Explanation texts for the local and central models.
@@
Figure 1. The final version of our privacy label for the local model. The central version is included in Appendix D.
@@
We use an 8 x 2 experimental design (all conditions listed in Appendix D), varying both the explanation type and the deployment model—local or central. We evaluate the privacy labels (Figure 1), process text, and metaphor text individually as well as in combination: metaphor + process, metaphor + process + label, metaphor + label, and process + label. We compare these seven conditions against the implication-focused explanations from (Xiong et al., 2020). Although (Xiong et al., 2020) evaluated a number of different explanations, we chose to compare against the explanation that led to the highest comprehension of privacy protections. Table 2 provides all three text explanations.
@@
Our goal is to set privacy expectations appropriately. Thus, we use a series of true/false questions from prior work about whether certain types of disclosure are possible to measure objective comprehension (Table 1). We also ask respondents about their subjective understanding of the explanations and how thorough they perceive the explanations to be (Kelley et al., 2009). Additionally, we ask whether respondents trust the non-profit organization to protect patient privacy (Xiong et al., 2020), and we ask two questions related to self-efficacy in decision making (Nanayakkara et al., 2023). All five questions use 5-pt Likert scales. Finally, although we do ask about willingness to share data with the non-profit, we caution against using this as a measure of explanation quality. The explanation that convinces the most people to share their data is not necessarily the best explanation. For example, we hope that a patient who is particularly concerned about disclosure to law enforcement would choose not to share data when it is protected using the central model.
@@
698 total respondents were recruited through Prolific, using the “balanced sample” feature—in accordance with best practices—to recruit an approximately representative sample in terms of gender (Tang et al., 2022). We conducted a power analysis to estimate an appropriate sample size; due to the large number of experimental conditions, we lack the statistical power to detect very small effects, but such small effects are unlikely to be meaningful in real-world contexts (Redmiles et al., 2018). Respondents were paid \$2 for completing the survey, and the median completion time was just under six minutes. A detailed breakdown of respondent demographics can be found in Appendix C.
@@
Table 3. Left: results from linear regression models for objective comprehension. We report regression coefficients ($\beta$) and 95% CIs for these coefficients. $\beta>0$ indicates an increase while $\beta<0$ indicates a decrease. Right: results from ordinal regression models for subjective understanding. We report odds ratios (OR) and corresponding 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease. For both columns, we use the Xiong et al. explanation as the reference level explanation.* p$<0.05$; ** p$<0.01$; *** p$<0.001$.
Analysis.
We analyze [^5] Analysis code: https://osf.io/3acvw/?view_only=f12174861ffd4cd0872a54a8e1326a26 the effect of our explanations on our dependent measures. We construct a set of regression models studying the effect of our independent variables—explanation and model—on our dependent variables: objective comprehension, subjective understanding, perceived thoroughness, trust, self-efficacy, and data-sharing decision. We use logistic regression models to study data-sharing decisions, linear regression models to study comprehension, and ordinal regression models to study the remaining dependent variables. In all models, we control for internet skill (Hargittai and Micheli, 2019).
@@
Table 4. Results from regression models for trust, perceived thoroughness, and self-efficacy, with the Xiong et al. explanation as the reference level explanation. Again we report odds ratios with 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease.
4.1. Results
4.1.1. Effectiveness of Designs
Comprehension.
[…]
Other Evaluation Criteria.
[…]
Feedback.
4.1.2. Prior Familiarity with PETs
@@
Table 5. Results from regression model for data-sharing decision. We report odds ratios (OR) and corresponding 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease.
4.1.3. Data-Sharing Decision
Respondents are more willing to share data (Table 5) under the local model ($\text{OR}=1.46$; $p<0.05$). This replicates findings from prior work and is likely due to the fact that the local model offers stronger privacy guarantees (Xiong et al., 2020). None of the explanations had a significant effect on data-sharing decisions.
When people are deciding whether to share information, they consider many other factors in addition to privacy protections (Smart et al., 2022; Naeini et al., 2017; Frik et al., 2023). In fact, many respondents simply were not worried about privacy (n=75). For example, one respondent felt that they had nothing in their medical history that they would “need to hide or be particularly private about.” Other respondents were interested in sharing their information, because they value helping others, mentioning benefits of data sharing (n=151). In the words of one respondent: “I do not have a problem with sharing my records if it will help someone.” On the other hand, respondents who were less willing to share their data often indicated that they felt it would be too risky or that their medical information was simply too private (n=242). For example, one respondent explained they were “not comfortable sharing [their] medical records with anyone but [their] doctor.” Other respondents wanted more information before they would be willing to share their data (n=155). However, the information they requested was not always related to DP. For example, some respondents wanted to know more about the non-profit organization. Finally, some participants distrusted either the non-profit or the privacy protection (n=88). In the words of one respondent:

### v2

@@ 3.1 Initial Prototypes
We also draw inspiration from prior work on visualizations of DP [13, 75, 82, 92, 78, 77] to design our own diagrams that highlight how DP protects or fails to protect against the disclosures listed in Table 1. We developed our diagrams through an iterative process. We discussed the accuracy and clarity of initial diagrams as a group, and based on the discussion, iterated on our designs. In the end, we developed four diagrams—two for the local model and two for the central model—with slight differences in iconography. After initial interviews, we added a third variation for both models that included a caption. All diagrams used a vertical line to depict the “privacy barrier,” as in [78], and used icons—most selected from the Noun Project [^3] https://thenounproject.com—to represent the different kinds of disclosures. Instead of using an illustration of a database, as in [78, 106], we use an icon of a filing cabinet to represent the collected data. Representative diagrams can be found in Appendix D.
Following guidance from prior work, we also developed privacy labels to clearly demonstrate which kinds of information disclosures DP can protect against. Each row corresponds to a specific information disclosure and clarifies whether protection is offered against said disclosure. We tested three different versions of the tables (six distinct tables in total, across the two models). One version of the table listed only the disclosures against which DP can protect. Thus for the local model, this table had five rows, whereas for the central model, this version had only one row. This table uses a red circle-backslash symbol to indicate that a particular disclosure is not permitted. The other two versions always included information about all five disclosures, but used different iconography to depict protection or lack thereof. Both of these versions incorporated lock icons to indicate when DP protected against a particular kind of disclosure. In one of these tables, we use a green lock icon—as recommended in prior work on connection security icons [32]—to indicate safety, whereas a red unlocked icon indicates disclosures against which DP does not protect. The other table is in black-and-white and uses the presence or absent of a lock icon to indicate (lack of) protection. Chrome previously used lock icons to indicate connection security, but has recently backed away from this choice due to concerns about overtrust; some Chrome users incorrectly assumed that a lock icon was a reflection on the safety of the website itself rather than the connection [16, 66]. Varying the use of icons allowed us to evaluate their appropriateness in a DP context. Appendix D includes representative versions of our original privacy labels.
3.2 Protocol
@@
Next, participants are informed that the non-profit plans to “use an extra layer of privacy protection in order to protect patients’ medical information.” Then, they are shown the first explanation of this privacy protection. After reading the explanation, the participants are asked to explain how patient data will be protected in their own words, as in [38]. Next, they are asked how they feel about the explanation, how well they feel that they understand the privacy protection after reading the explanation, what concerns they would have about sharing their data, and what else they would like to know about how patient data will be protected, adapting questions from [86]. If the design under discussion includes the use of color, they are also asked about these color choices. Finally, they are asked how the explanation could be improved.
@@ 3.3 Participant Recruitment
3.4 Analysis
The interviewer first transcribed and summarized all the interviews. Next, an interview from each condition was selected at random, forming a set of six interviews. The first two authors reviewed these six transcripts to develop a set of codes, organized into four distinct themes (Appendix A). They then shared this codebook with the research team and modified it based on the group’s feedback. Finally, the same two authors coded all 24 interviews together [^4] We do not calculate or report inter-rater reliability (IRR) for two reasons. One, while calculating IRR can be useful to establish agreement before researchers divide a corpus to code different subsets individually, in our case both researchers coded all of the data together. Two, we are not seeking to make quantitative claims about our codes [68]. with the updated codebook.
3.5 Findings
3.5.1 Effectiveness of Initial Designs
@@
Responses to the metaphors were mixed. Some participants appreciated the concision of the metaphors, while others wanted more details. For example, one participant criticized an explanation’s brevity, saying it is “a little bit simple and […] doesn’t go into too many details.” (P8) In contrast, a different participant complimented this very quality by describing an explanation as “reader-friendly, very concise” (P5). This tension between accuracy and thoroughness of explanations on the one hand, and simplicity on the other has also been reported in other domains, such as explainable machine learning [1] and privacy policies [37].
@@
Despite the fact that the diagrams showed the hacker to the left of the privacy barrier, two of the four participants in this condition nevertheless explicitly stated that the privacy barrier would protect their data from hackers. Many people may expect PETs to protect against hackers and criminals, making the protection offered by central DP alone somewhat unintuitive [92]. We dropped the diagram explanations due to the pervasive confusion expressed by participants.
Xiong et al. [106] previously investigated the use of diagrams for explaining location privacy and found less than ideal levels of comprehension, particularly for the local model, though they speculate that data quality issues with Amazon Mechanical Turk may be to blame. Alternatively, it is possible that data flow diagrams inherently overemphasize processes at the expense of clearly enumerating implications.
@@
All of our explanations were designed to communicate the implications of DP rather than the details of how DP works. Prior work has shown that explaining the process of adding noise to data is not enough to help people understand the consequences data sharing [105]. Nevertheless, omitting any discussion of process seems to leave people unsatisfied and confused. Most participants had questions about how the data protection worked. Providing a detailed explanation of the mathematical and technical details of DP is likely to overwhelm most people, but people nevertheless do want some information about how DP works—finding the right balance may be challenging. This finding aligns with prior work on explaining encryption. While explanations of encryption focused on outcome lead to greater perceived security than explanations focused on process, hybrid explanations that incorporate information on both process and outcome lead to the greatest perceived security [26]. Prior work on metaphors for DP also found that some participants were interested in understanding how DP works [51].
@@
[skipabove=] Design Changes: Another text was added to provide context about how DP works; we adapted a state-of-the-art text explanation by Xiong et al. [105], while aiming for improved readability by eliminating terms like “database” and “aggregated.” We anticipated that this additional information on process could complement our other explanations that focus on implications.
3.5.2 Mental Models
@@
Some participants—especially in the diagram and privacy label conditions—reasoned about DP through comparisons to other PETs. For example, one participant understood the privacy barrier as “some kind of firewall that keeps [their] privacy safe” (P21). Encryption in particular was mentioned frequently, perhaps because it is a particularly familiar and ubiquitous PET or perhaps because participants associated our lock icons with encryption [32, 42]. One participant, who assumed that encryption was the technology being described, wanted to know “what type of encryption” (P23) was used. Prior work has also found associations between DP and encryption, and found that associations with encryption correspond to higher trust [51]. DP is distinct from encryption, so while it may be possible to leverage people’s knowledge about encryption to construct better explanations of DP, it is also likely that associations with encryption may lead to misconceptions.
@@
Once again, while this model oversimplifies DP, it shares key elements with the truth and thus is likely useful overall. However, it is important for people to understand that the “fake” data nonetheless reveals useful information about the overall distribution; therefore, DP does not necessarily offer protection against inferential privacy risks [55, 56].
3.5.3 Validating Design Changes.
@@
4 Large-Scale Evaluation
@@
Figure 1: The final version of our privacy label for the local model. The central version is included in Appendix D.
@@
Respondents who pass the comprehension check are then asked whether they trust the non-profit to protect patient privacy [105]. Next they are asked two questions related to self-efficacy. Finally, they are asked whether they would be willing to share their information with the non-profit. An open-text box asks them to explain their decision.
Respondents then answer five true/false questions on privacy expectations, followed by the Likert-scale questions about understanding and thoroughness. They are also invited to share feedback on the explanations in a free-response text box. When answering the above questions, respondents have the option to reread the descriptions of the scenario and privacy protection at any time. Next, respondents are asked about their familiarity with various PETs, including DP and a non-existent technology (“deliquescent security”). If they indicate familiarity with some of the listed technologies, they are asked which of the technologies (if any) was described in the survey. In a free-response text box, they are asked to explain their reasoning. Finally, respondents answer questions about themselves. In addition to standard demographic questions (i.e., age, income, race, ethnicity, gender, education, job field), the survey also includes measures of internet skill [40]. The full survey instrument is included in Appendix B.
Table 2: Explanation texts for the local and central models.
@@
We use an 8 x 2 experimental design (all conditions listed in Appendix D), varying both the explanation type and the deployment model—local or central. We evaluate the privacy labels (Figure 1), process text, and metaphor text individually as well as in combination: metaphor + process, metaphor + process + label, metaphor + label, and process + label. We compare these seven conditions against the implication-focused explanations from [105]. Although [105] evaluated a number of different explanations, we chose to compare against the explanation that led to the highest comprehension of privacy protections. Table 2 provides all three text explanations.
@@
Our goal is to set privacy expectations appropriately. Thus, we use a series of true/false questions from prior work about whether certain types of disclosure are possible to measure objective comprehension (Table 1). We also ask respondents about their subjective understanding of the explanations and how thorough they perceive the explanations to be [53]. Additionally, we ask whether respondents trust the non-profit organization to protect patient privacy [105], and we ask two questions related to self-efficacy in decision making [77]. All five questions use 5-pt Likert scales. Finally, although we do ask about willingness to share data with the non-profit, we caution against using this as a measure of explanation quality. The explanation that convinces the most people to share their data is not necessarily the best explanation. For example, we hope that a patient who is particularly concerned about disclosure to law enforcement would choose not to share data when it is protected using the central model.
@@
698 total respondents were recruited through Prolific, using the “balanced sample” feature—in accordance with best practices—to recruit an approximately representative sample in terms of gender [96]. We conducted a power analysis to estimate an appropriate sample size; due to the large number of experimental conditions, we lack the statistical power to detect very small effects, but such small effects are unlikely to be meaningful in real-world contexts [87]. Respondents were paid \$2 for completing the survey, and the median completion time was just under six minutes. A detailed breakdown of respondent demographics can be found in Appendix C.
Analysis.
We analyze [^5] Analysis code: https://osf.io/3acvw/?view_only=f12174861ffd4cd0872a54a8e1326a26 the effect of our explanations on our dependent measures. We construct a set of regression models studying the effect of our independent variables—explanation and model—on our dependent variables: objective comprehension, subjective understanding, perceived thoroughness, trust, self-efficacy, and data-sharing decision. We use logistic regression models to study data-sharing decisions, linear regression models to study comprehension, and ordinal regression models to study the remaining dependent variables. In all models, we control for internet skill [41].
@@
Table 3: Left: results from linear regression models for objective comprehension. We report regression coefficients ($\beta$) and 95% CIs for these coefficients. $\beta>0$ indicates an increase while $\beta<0$ indicates a decrease. Right: results from ordinal regression models for subjective understanding. We report odds ratios (OR) and corresponding 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease. For both columns, we use the Xiong et al. explanation as the reference level explanation.* p$<0.05$; ** p$<0.01$; *** p$<0.001$.
4.1 Results
4.1.1 Effectiveness of Designs
Comprehension.
[…]
@@ 4.1.1 Effectiveness of Designs
Table 4: Results from regression models for trust, perceived thoroughness, and self-efficacy, with the Xiong et al. explanation as the reference level explanation. Again we report odds ratios with 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease.
@@
Table 5: Results from regression model for data-sharing decision. We report odds ratios (OR) and corresponding 95% CIs. An OR $>1$ indicates an increase in odds, while an OR $<1$ indicates a decrease.
Other Evaluation Criteria.
[…]
Feedback.
4.1.2 Prior Familiarity with PETs
4.1.3 Data-Sharing Decision
Respondents are more willing to share data (Table 5) under the local model ($\text{OR}=1.46$; $p<0.05$). This replicates findings from prior work and is likely due to the fact that the local model offers stronger privacy guarantees [105]. None of the explanations had a significant effect on data-sharing decisions.
When people are deciding whether to share information, they consider many other factors in addition to privacy protections [92, 74, 35]. In fact, many respondents simply were not worried about privacy (n=75). For example, one respondent felt that they had nothing in their medical history that they would “need to hide or be particularly private about.” Other respondents were interested in sharing their information, because they value helping others, mentioning benefits of data sharing (n=151). In the words of one respondent: “I do not have a problem with sharing my records if it will help someone.” On the other hand, respondents who were less willing to share their data often indicated that they felt it would be too risky or that their medical information was simply too private (n=242). For example, one respondent explained they were “not comfortable sharing [their] medical records with anyone but [their] doctor.” Other respondents wanted more information before they would be willing to share their data (n=155). However, the information they requested was not always related to DP. For example, some respondents wanted to know more about the non-profit organization. Finally, some participants distrusted either the non-profit or the privacy protection (n=88). In the words of one respondent: